Bob uses p = (141, 19) as his public key and S = 21 as his secret key. Is Bob's system correct? Please show all work and explanations. Thank You
Answer #1

p=(141,19) that means N=141=3×47

r=(3-1)(47-1)=92

We choose k such that k=1 mod r

Let k=93, factors of k=3×31

we have e = 19
d = 21
N = 141
r = 92
e*d = 399
e*d mod r = 31
e and r are relatively prime
d and r are relatively prime. This implies e and d are consistent.

This gives us that the public key and the secret key are acceptable.

Hence, Bob's system is correct.